OpenStudy (anonymous):

If f(x)=(1+x)/x, for which value of x does f(x)=4

jimthompson5910 (jim_thompson5910):

Replace f(x) with 4 and solve for x f(x)=(1+x)/x 4=(1+x)/x 4x = 1+x ... ... ... x = ???

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

what did you get for x?

OpenStudy (anonymous):

1/3?

OpenStudy (anonymous):

looks good to me

OpenStudy (anonymous):

thanks